OpenStudy (natsent06):

NEED ANSWER FAST!!!! You treated yourself and a friend to sodas and snacks at a concert. The total was $16.75. You tipped 15%, and there was a sales tax of 6.5%. What was the total amount you paid?

OpenStudy (anonymous):

16.75 * 0.15 = a 16.75 * 0.65 = b 16.75 + a + b

OpenStudy (anonymous):

I mean 16.75 * 0.065 = b

OpenStudy (natsent06):

is it $20.35?

OpenStudy (anonymous):

Yes.

OpenStudy (natsent06):

thanks!